33.79 mL of a 0.157 M solution of weak acid HA are titrated with 0.22 M NaOH. What is the pH of the solution after 12.98 mL of the NaOH have been added? The Ka for the weak acid is 0.0000005732.
Answer #1

no of moles of HA    = molarity*volume in L

                                 = 0.157*0.03379

                                  = 0.00530503 moles

no of moles of NaOH   = molarity*volume in L

                                     = 0.22*0.01298   = 0.0028556 moles

PKa   = -logKa

          = -log 0.0000005732

          = 6.2417

                HA(aq)     +     NaOH(aq) -------------------> NaA(aq) + H2O(l)

I                0.00530503     0.0028556                         0

C              -0.0028556      -0.0028556                         0.0028556

E             0.00244943           0                                    0.0028556

            PH    = PKa + log[NaA]/[HA]

                      = 6.2417 + log0.0028556/0.00244943

                      = 6.2417 + 0.06663   = 6.30833